What is 1/3 of 6?

Found 3 solutions by usyim88hk, sschoentrup, lenny460:
Answer by usyim88hk(157)

It is 6 times 1 and then divide by 3
which is (6*1)/3 = 6/3 = 2

Answer by sschoentrup(1)


1/3 x 6/1 = 6/3 = 2/1 = 2

Answer by lenny460(775)

1/3 * 6/1
= 6/3
= 2/1
= 2